OpenStudy (anonymous):

Calculus help? Please? Find the equation for y, given the derivative and the indicated point on the curve. dy/dx = 3x^2-5; y(0)=1

ganeshie8 (ganeshie8):

take integral both sides

ganeshie8 (ganeshie8):

dy/dx = 3x^2-5 y = integral (3x^2-5) dx

OpenStudy (anonymous):

How do I integrate that one? do I use x^n dx = x^(n+1)/(n+1) +C?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@mathmale ?

OpenStudy (mathmale):

Hello, VBG!! I view this as a separable differential equation. You comfortable with that? If so, please separate x-terms from y-terms and write out the result.

OpenStudy (mathmale):

BRB

OpenStudy (anonymous):

If thats what you think I agree. I learn from you guys! :) Ok so I should solve 3x^2 and -5 separately?

OpenStudy (anonymous):

integrate I mean

OpenStudy (mathmale):

Let me type that out in Equation Editor for added clarity:

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

\[\frac{ dy }{ dx}=3x ^{?}-5\]

OpenStudy (mathmale):

becomes \[dy=(3x ^{2}-5)dx\]

OpenStudy (mathmale):

Why is this advantageous?

OpenStudy (anonymous):

advantageous? because it is easier to solve?

OpenStudy (mathmale):

Yes, VBG, you do need to integrate 3x^2 and -5 separately. Add a constant of integration as well.

OpenStudy (mathmale):

I'd prefer you explain that what I did is advantageous because it emphasizes that yours is a separable differential equation.

OpenStudy (mathmale):

\[\int\limits_{}^{}dy = \int\limits_{}^{}(3x ^{2}-5)dx\]

OpenStudy (anonymous):

right ok. so integral 3x^2 = x^3 + C

OpenStudy (anonymous):

and integral -5 = -5x + C

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

mathamle I have to go and will be back in an hour or so!! hopefully you will be on :)

OpenStudy (anonymous):

@mathmale Hi sorry for taking so long!!

OpenStudy (mathmale):

The right side of your integral now becomes x^3-5x + C. What's the left side? The initial condition says y(0)=1. In y=x^3-5x + c, let x=0 and y=1. Find C. Then write the solution in the form y=x^3-5x + C.

OpenStudy (anonymous):

the left side would just show the integral x^n dx? Ok so y= 0^3 - 5(0) + C So C= 1

OpenStudy (anonymous):

yes?

OpenStudy (mathmale):

Then y=x^3-5x + 1. Check this by substitution of y=x^3 - 5x + 1 and y ' = 3x^2 -5 back into the original equation. Sorry for the delay; I messed up a display setting on my computer and had to re-learn how to correct that. Hope to see you again soon on OpenStudy!

OpenStudy (anonymous):

yep thats what I did. No problem thanks!